OpenStudy (anonymous):

Select the equation of the line that passes through the point (–2, 1) and has slope 3 in point-slope form.

OpenStudy (anonymous):

y-y1 = m(x-x1) y-1 = 3 (x+2) y-1 = 3x+6 y=3x+7

OpenStudy (anonymous):

(y-1)/(x+2) = 3

OpenStudy (amistre64):

...shudder... point slope format; ewww!!

OpenStudy (amistre64):

a(x-x1) +b(y-y1) = 0

OpenStudy (amistre64):

<a,b> is normal to the plane :)